# ESC-Eval: Evaluating Emotion Support Conversations in Large Language Models

### v2

@@ Abstract
Emotion Support Conversation (ESC) is a crucial application, which aims to reduce human stress, offer emotional guidance, and ultimately enhance human mental and physical well-being. With the advancement of Large Language Models (LLMs), many researchers have employed LLMs as the ESC models. However, the evaluation of these LLM-based ESCs remains uncertain. Inspired by the awesome development of role-playing agents, we propose an ESC Evaluation framework (i.e., ESC-Eval), which uses a role-playing agent to interact with ESC models, followed by a manual evaluation of the interactive dialogues. In detail, we first re-organize 2,801 role-playing cards from seven existing datasets to define the roles of the role-playing agent. Second, we train a specific role-playing model - ESC-Role to mimic the behavior of a real person experiencing distress. Third, through ESC-Role and organized role cards, we systematically conduct experiments using 14 LLMs as the ESC models, including general AI-assistant LLMs (e.g., ChatGPT) and ESC-oriented LLMs (e.g., ExTES-Llama). We conduct comprehensive human annotations on interactive multi-turn dialogues of different ESC models. The results show that ESC-oriented LLMs exhibit superior ESC abilities compared to general AI-assistant LLMs, but there is still a gap behind human performance. Moreover, to automate the evaluation of future ESC models, we developed ESC-RANK, which trained on the annotated data, achieving a scoring performance surpassing 35 points of GPT-4. Our data and code are available at https://github.com/haidequanbu/ESC-Eval.
@@
Haiquan Zhaothanks: Work done during internship at Shanghai Artificial Intellignece Laboratory1,2   Lingyu Li2   Shisong Chen1   Shuqi Kong2    Jiaan Wang 1 Kexin Huang1,2   Tianle Gu2   Yixu Wang1,2   Jian Wang2   Dandan Liang2 Zhixu Lithanks: Correspondence to: Zhixu Li<zhixuli@fudan.edu.cn> and Yan Teng <tengyan@pjlab.org.cn>1  Yan Teng22footnotemark: 22    Yanghua Xiao1   Yingchun Wang2 1 Fudan University 2 Shanghai Artificial Intelligence Laboratory
@@ 3.3 Correlation Analysis
Dimen. Model ACC ACCsoft
Flue. InternLM2 31.84/17.15 93.07/72.82
@@ 5 Conclusion
The crowdsourced annotators in this article are not native English speakers, but all of them are proficient English users, and all of them are master’s students or PhD candidates in the humanities and social sciences. However, they still cannot avoid possible shortcomings in English annotation.
@@ A.1 Resource Datasets
Figure 4: The framework of user-card construction. Firstly, the initial user cards are extracted from open-source datasets using GPT-4. In the second step, based on the scene classification we designed, GPT-4 is utilized to determine the category to which the character sheet data belongs, and further filtering is performed. In the third step, we employ crowdsourcing to annotate the headphone category and subcategories of the scenes, and manually filter the user cards again.
@@ A.3 Data Analysis
Table 7: The quantities of role cards at each stage. Extract represents the initial number of role cards extracted from open resource datasets; Filter1 represents the number of role cards after the filtering process using GPT-4; Filter2 represents the number of role cards after manual filtering; High represents high-quality role cards, and Middle represents medium-quality role cards.
@@ C.3 Correlation Analysis
Skillful Focused on five aspects: 1. Empathy 2. Information 3. Hopeful 4. Importance 5. Providing necessary advice, or highlighting bright spots. One out of five. 0

### v3

@@ Abstract
Emotion Support Conversation (ESC) is a crucial application, which aims to reduce human stress, offer emotional guidance, and ultimately enhance human mental and physical well-being. With the advancement of Large Language Models (LLMs), many researchers have employed LLMs as the ESC models. However, the evaluation of these LLM-based ESCs remains uncertain. Inspired by the awesome development of role-playing agents, we propose an ESC Evaluation framework (i.e., ESC-Eval), which uses a role-playing agent to interact with ESC models, followed by a manual evaluation of the interactive dialogues. In detail, we first re-organize 2,801 role-playing cards from seven existing datasets to define the roles of the role-playing agent. Second, we train a specific role-playing model - ESC-Role to mimic the behavior of a real person experiencing distress. Third, through ESC-Role and organized role cards, we systematically conduct experiments using 14 LLMs as the ESC models, including general AI-assistant LLMs (e.g., ChatGPT) and ESC-oriented LLMs (e.g., ExTES-Llama). We conduct comprehensive human annotations on interactive multi-turn dialogues of different ESC models. The results show that ESC-oriented LLMs exhibit superior ESC abilities compared to general AI-assistant LLMs, but there is still a gap behind human performance. Moreover, to automate the evaluation of future ESC models, we developed ESC-RANK, which trained on the annotated data, achieving a scoring performance surpassing 35 points of GPT-4. Our data and code are available at https://github.com/AIFlames/Esc-Eval.
@@
Haiquan Zhaothanks: Work done during internship at Shanghai Artificial Intellignece Laboratory1,2   Lingyu Li1   Shisong Chen2   Shuqi Kong1    Jiaan Wang 2 Kexin Huang1,2   Tianle Gu1   Yixu Wang1,2   Jian Wang1   Dandan Liang1 Zhixu Lithanks: Correspondence to: Zhixu Li<zhixuli@ruc.edu.cn> and Yan Teng <tengyan@pjlab.org.cn>3,4  Yan Teng22footnotemark: 21    Yanghua Xiao2   Yingchun Wang1 1 Shanghai Artificial Intelligence Laboratory 2 School of Computer Science, Fudan University 3 School of Information, Renmin University of China 4 Suzhou Key Laboratory of Artificial Intelligence and Social Governance Technologies, International College (Suzhou Research Institute), Renmin University of China
@@ 3.3 Correlation Analysis
Dim. Model ACC ACCsoft
Flu. InternLM2 31.84/17.15 93.07/72.82
@@ 5 Conclusion
The crowdsourced annotators in this article are not native English speakers, but all of them are proficient English users. We employed a total of 14 annotators. Among them, one has a background in computer science, two in mental health sciences/psychology, and the remaining 11 in humanities and social sciences such as law, sociology, and history. Three are PhD candidates in China, while the others are master’s students. However, they still cannot avoid possible shortcomings in English annotation.
The role cards published in this paper were extracted using GPT4. As much as possible this project uses a variety of filtering methods, including model filtering and manual filtering, there is still no guarantee of potential bias in the models included in the role cards. We hope future projects can improve.
@@
Acknowledgements
This work was supported by the National Key R&D Program of China (2022ZD0160103), Shanghai Science and Technology Innovation Action Plan (No. 22511104700) and Shanghai Artificial Intelligence Laboratory.
@@ A.1 Resource Datasets
Figure 4: The framework of user-card construction. Firstly, the initial user cards are extracted from open-source datasets using GPT-4. In the second step, based on the scene classification we designed, GPT-4 is utilized to determine the category to which the character sheet data belongs, and further filtering is performed. In the third step, we employ crowdsourcing to annotate the category and subcategories of the scenes, and manually filter the user cards again.
@@ A.3 Data Analysis
Table 7: Some cases of different quality role cards.
@@ B.3 Findings
Issues when using ESC-Role.
ESC-Role is trained on Qwen1.5-14B-Chat, which could lead to certain issues like preference bias when interacting with models based on Qwen. Furthermore, when conversing with general and domain ESC models, ESC-Role might generate diverse responses due to the different lengths of inputs of ESC models under evaluation. However, this variability in responses aligns more closely with real user scenarios.
@@ C.3 Correlation Analysis
Skillful 1. Empathy 2. Information 3. Hopeful 4. Importance 5. Providing necessary advice, or bright spots. One out of five. 0